OpenStudy (anonymous):

quick question... is the reciprocal of a number always 1/the number? for example, the reciprocal of 2 would be 1/2?

OpenStudy (anonymous):

Yes except 0.

OpenStudy (anonymous):

same with fractions?

OpenStudy (anonymous):

Yes

jimthompson5910 (jim_thompson5910):

reciprocal of 1/3 is 1/( 1/3 ) = 3/1 = 3

jimthompson5910 (jim_thompson5910):

so you're just flipping the fraction

OpenStudy (anonymous):

thank you